​Eagletron's current stock price is $ $10. Suppose that over the current​ year, the stock price will either increase by 96% or decrease by 51%. ​Also, the​ risk-free rate is 25% ​(EAR).

a. What is the value today of a​ one-year at-the-money European put option on Eagletron​ stock?

b. What is the value today of a​ one-year European put option on Eagletron stock with a strike price of $19.60​?

c. Suppose the put options in parts ​(a​) and ​(b​) could either be exercised​ immediately, or in one year. What would their values be in this​ case?
Answer #1
Answer a)
Pu= 0 Pd= 5.1
delta= (0-5.1)/(19.6-5.1) -0.35172
B= (5.1-5.1(-0.35172))/1.25 5.515018
Price today= Delta* stock price +B
1.998
Answer b)
The payoff are Pu = 0 pd= 9.6+4.9 =14.5
so 14.5/5.1 2.843137255
value should be 5.6800
Answer c)
Intrinsic value in a)= 0 so we should exercise imediately as we are getting $1.998
Intrinsic value in b)= 19.6-10=9.6 so we should exercise now